Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 370     Entry time: 27 Sep 2023, 07:51
Author: Benjamin van Kuiken 
Group: SCS 
Subject: SCS Night Shift Summary 

Achievements

  • Characterized optical laser spot at sample

Problems

  • FEL beam had very unstable intensity and spectrum. As a result no useful data could be collected during the night
    • FEL pointing seemed to be stable on all tunnel imagers. The issue was communicated to BKR, and they said that things looked fine from there perspective.
    • Large fluctuations in spectral distrubtion at SCS spectrometer were observed. See video at https://syncandshare.xfel.eu/index.php/s/emtwonCsQJJqpz6
    • Oth-order beam on SCS pop-in showed large intensity fluctuations with ~50% of the time showing no beam at all
Attachment 1: SCSPopinOthOrder_270923.png  255 kB  | Hide | Hide all
SCSPopinOthOrder_270923.png
ELOG V3.1.4-7c3fd00